OpenStudy (anonymous):

√x = 2/5 Write final answer as a fraction with no spaces.

OpenStudy (anonymous):

square both sides

OpenStudy (anonymous):

\[x = (\frac{2}{5})^{2} = \frac{4}{25}\]

OpenStudy (saifoo.khan):

\[\sqrt x = \frac25\]Taking Square on both sides,\[(\sqrt x)^2 = (\frac25)^2\]\[x = \frac{2^2}{5^2}\] \[x = \frac{4}{25}\]

OpenStudy (anonymous):

Thank you SOOO MUCH!